Suppose that each firm currently catches $100 worth of fish per year, and there is $200 worth of fish left in the lake. In the following spreadsheet, enter the amount that each firm would earn in each of the next 10 years if the two firms continue to fish.


1
Expert's answer
2022-02-16T16:49:16-0500

Question is incomplete. Data is missing for fish catching rate by each firm.

Assume fish catching rate by one firm is 2% each year.


Year Amount

1 100+100(2%)=102

2 102+102(2%)=104.04

3 104.04+104.04(2%)=106.1208

4 106.1208+106.1208(2%)=108.243216

And so on using the formula "100(1+2\%)^n" , where n is the number of year.